OpenStudy (anonymous):

2 2/5 divided by 1 3/4 = 2 3/4 divided by 4 1/2 =

Parth (parthkohli):

5/2 / 7/4 = 5/2 * 4/7 = 20/14 = 10/7 = 1 3/7

Parth (parthkohli):

11/4 / 9/2 = 11/4 * 2/9 = 22/36 = 11/18